# Harry Yang1, Linfeng Zhang3, Qifeng Chen1🖂, Yue Ma1🖂

### v2

@@ Abstract
footnotetext: † Equal contribution.footnotetext: 🖂 Corresponding author.
@@ 1 Introduction
Image editing methods based on diffusion [5, 45, 10] and flow models [20, 18, 19, 32, 17] have demonstrated considerable success in general tasks, yet they often fail when faced with complex, large-scale shape transformations. These models can struggle to modify an object’s structure as intended or may inadvertently alter background regions, which degrades the overall image quality. This limitation indicates a critical gap in their ability to perform precise structural edits while maintaining the integrity of unedited content. Meanwhile, the rise of AIGC foundation models has significantly accelerated progress across various downstream applications, including visual generation and editing [31, 33, 35, 29, 30, 50, 28, 6, 11], animation and colorization [34, 57].
@@
To address these challenges, we propose Follow-Your-Shape, a training-free framework for precise and controllable shape editing. As illustrated in Figure 2, the core innovation of our pipeline is the Trajectory Divergence Map (TDM). The TDM is generated by computing the token-wise difference between the denoising velocity fields of the source and target prompts. This map accurately localizes the regions intended for editing, which in turn guides a selective KV injection mechanism to ensure that modifications are applied precisely where needed while preserving the background.
However, directly applying TDM-guided injection across all denoising timesteps is suboptimal because the TDM can be unstable in the early, high-noise stages of the process. We therefore introduce a Scheduled KV Injection strategy that adapts its guidance throughout the denoising process. As visualized in Figure 2, this staged approach first performs unconditional KV injection to stabilize the initial trajectory, and only then applies TDM-guided editing once a coherent latent structure has formed. This scheduling ensures a more robust and faithful editing outcome compared to a direct application.
To validate our approach, we introduce ReShapeBench, a new benchmark for the rigorous evaluation of large-scale shape modifications. On this benchmark, Follow-Your-Shape demonstrates state-of-the-art performance. Quantitative results show that our method achieves superior background preservation, with a PSNR [14] of 35.79 and an LPIPS [56] score of 8.23. Furthermore, our method attains the highest scores for both text-to-image alignment (CLIP-Sim [40]: 33.71) and overall aesthetic quality, measured using the LAION aesthetic predictor [43] (Aesthetic Score: 6.57), confirming its effectiveness.
@@
  • A novel and training-free editing framework, Follow-Your-Shape, that uses a Trajectory Divergence Map (TDM) to achieve precise, large-scale shape transformations while preserving background content.
@@ 2 Related Work
2.1 Region-Specific Image Editing
A central challenge in image editing is localizing modifications to specific regions [4, 60, 58, 21, 32, 13]. Early methods often relied on explicit user-provided masks to delineate editable areas [24, 1, 7, 53, 49]. While effective for certain tasks, this approach requires manual annotation, limiting its applicability. To address this, subsequent work explored techniques to infer editable regions directly from text prompts. Methods such as Prompt-to-Prompt [12] and Plug-and-Play [45] manipulate cross-attention maps to associate textual tokens with spatial areas, enabling localized edits without explicit masks. Other approaches, such as DiffEdit [9], generate a mask by computing differences between diffusion model predictions conditioned on source and target prompts. However, attention-based localization can be imprecise and unstable, especially during large-scale shape transformations where object boundaries change significantly [37, 5]. In contrast, Follow-Your-Shape provides a training-free and mask-free method for identifying editable regions directly from the model’s behavior, avoiding the need for external masks or noisy attention maps.
2.2 Structure Preservation via Inversion and Feature Reuse
Preserving non-target regions is equally critical for high-fidelity editing, and this is closely tied to the quality of the model’s inversion process. For diffusion models, significant research has focused on improving DDIM inversion [44] to better reconstruct a source image from noise. Previous works like null-text inversion [36] and optimization-based methods [47] aim to reduce the discrepancy between the reconstruction and editing trajectories. With the shift toward flow-based models, inversion fidelity has become even more important due to their deterministic nature. Methods like RF-Solver [51] achieve more accurate reconstructions by incorporating higher-order derivative information. Beyond improving inversion, another line of work focuses on explicitly reusing modules or features from the source image’s generation process [59, 27, 54]. Techniques based on Key-Value (KV) caching [61, 2] or feature injection [51, 10] enforce structural consistency by propagating source-image features into the new generation process. In contrast to prior methods that rely on simple heuristics, Follow-Your-Shape employs a trajectory-guided scheduled injection strategy to achieve more precise, content-aware control.
@@ 3.1 Motivation
Effective image editing requires a precise balance between introducing new content and preserving the original structure. As illustrated in Figure 2 (Top), traditional structure-preserving editing approaches often produce unstable denoising trajectories that deviate significantly from the stable reconstruction path, leading to severe structural degradation and undesired artifacts. Moreover, prior methods for localizing edits have notable drawbacks:
  • Binary Segmentation Masks: Rely on external tools, introducing overhead and a dependency on mask quality. Their rigid boundaries hinder large-scale shape changes and often produce artifacts.
  • Cross-Attention Masks: Inferred from model attention, these maps are often noisy and inconsistent, proving unreliable for localizing edits, especially during significant shape transformations.
@@
Figure 2: Motivation for Trajectory Divergence Map (TDM) Guided Editing. Top: Standard editing methods (red) often produce unstable trajectories compared to the stable reconstruction path (orange), leading to distorted outputs. Middle: Our staged approach first stabilizes the editing trajectory before using the TDM to guide it toward the target concept. This method supports diverse shape modifications (dashed lines). Bottom: The TDM visualizes the dynamically localized editing region across different timesteps, with different border colors correspond to different stages.
Figure 3: Overview of our proposed pipeline. Given a source image and a corresponding prompt, we first perform inversion to obtain the initial noisy latent code $x_{T}$. The editing process is then divided into three stages. In Stage 1, we stabilize the initial denoising trajectory. This is achieved by injecting key-value (KV) features from the inversion path into the denoising model during its initial steps. In Stage 2, we compute a Trajectory Divergence Map (TDM) by comparing the denoising trajectories generated from the source and edit prompts. This map is then processed to precisely identify the regions intended for editing. In Stage 3, we perform the final edit. Guided by the TDM, blended KV features are injected into the final attention blocks of the denoising model to introduce the new semantics. Simultaneously, ControlNet conditions are supplied to ensure the edited regions conform to the original structure.
@@ 3.2 Follow-Your-Shape
Our objective is to perform shape-aware edits by selectively preserving background and non-target regions. We achieve this through a staged editing process that combines scheduled Key-Value (KV) injection with structural guidance, where the edit is localized by the Trajectory Divergence Map (TDM).
@@ 3.2.1 Trajectory Divergence Map
Our approach is grounded in the perspective of flow trajectories within the latent space, extending concepts from flow-matching frameworks to the inference setting. As illustrated in Figure 2 (Top), a standard reconstruction follows a stable denoising trajectory guided by the source prompt $\mathbf{c}_{\text{src}}$. In an editing task, conditioning on a target prompt $\mathbf{c}_{\text{tgt}}$ alters the velocity field, causing the denoising trajectory to deviate from this initial path. We posit that the magnitude of this deviation spatially localizes the semantic difference between the two prompts. Regions intended for modification will exhibit significant divergence, while background areas will follow nearly identical trajectories. To formalize this, let $\{\mathbf{x}_{t}\}_{t=0}^{T}$ be the latent sequence from the source image inversion, and let $\{\mathbf{z}_{t}\}_{t=0}^{T}$ be the corresponding sequence during the editing (denoising) process. We define the token-wise Trajectory Divergence Map (TDM) $\delta_{t}$ at timestep $t$ as the $L_{2}$ norm of the difference between the velocity vectors predicted under the two prompts:
@@
As shown in the bottom panel of Figure 2, this produces a normalized TDM, $\{\tilde{\delta}_{t}^{(i)}\}$, which quantifies the localized editing strength on a scale of $[0,1]$.
3.2.2 Scheduled KV Injection and Mask Generation
Directly applying TDM-guided injection across all timesteps is suboptimal due to the instability of the TDM in early, high-noise regimes. To address this, we introduce a scheduled injection strategy that partitions the $N$ denoising steps into three distinct phases, adapting the guidance mechanism to the state of the latent variable.
@@ 3.2.2 Scheduled KV Injection and Mask Generation
For an initial set of $k_{\text{front}}$ timesteps, we perform unconditional KV injection from the source inversion path across all spatial tokens. This operation enforces a global reconstruction objective, equivalent to setting the edit mask $M_{S}=\mathbf{0}$, which stabilizes the trajectory and prevents semantic drift while the latent representation $\mathbf{z}_{t}$ is still dominated by noise.
Stage 2: TDM-Guided Injection.
Once a stable latent structure has emerged, for the subsequent timesteps ahead of $k_{\text{tail}}$, we activate TDM-guided injection. To form a temporally consistent and spatially coherent edit mask, we first aggregate the normalized per-step TDMs $\{\tilde{\delta}_{t}\}$ from Eq. 2 over an editing window $T$. This is achieved via a softmax-weighted temporal fusion for each token $i$:
$\hat{\delta}^{(i)}=\sum_{t\in T}\alpha_{t}^{(i)}\cdot\tilde{\delta}_{t}^{(i)},\quad\text{where}\quad\alpha_{t}^{(i)}=\frac{\exp(\tilde{\delta}_{t}^{(i)})}{\sum_{t^{\prime}\in T}\exp(\tilde{\delta}_{t^{\prime}}^{(i)})}.$ (3)
The resulting map $\hat{\delta}$ is further refined via convolution with a Gaussian kernel $\mathcal{G}_{\sigma}$ to produce the final edit mask $M_{S}\in[0,1]^{H\times W}$:
$M_{S}=\mathcal{G}_{\sigma}*\hat{\delta}$. (4)
The smoothed attention map is then binarized with a threshold of $\tau=0.35$ to produce the final mask, $M_{S}$. This binary mask enables a selective fusion of Key-Value (KV) features. For regions where the mask is active ($M_{S}=1$), we utilize the target features ($K_{\text{tgt}}$, $V_{\text{tgt}}$) computed in the current denoising step. Conversely, for inactive regions ($M_{S}=0$), we inject the corresponding source features ($K^{\text{inv}}$, $V^{\text{inv}}$) that were stored during the inversion process. This feature-blending operation is formulated as:
$V^{*}\leftarrow M_{S}\odot V_{\text{tgt}}+(1-M_{S})\odot V^{\text{inv}}.$ (5)
An analogous operation is performed for the key feature $K$ as demonstrated in Algorithm 1.
@@
Our framework ensures editing conformance by combining explicit structural guidance via ControlNet with semantic preservation through our TDM-guided feature injection, which extends the principles of RF-Edit. To enforce strong geometric constraints, ControlNet conditions the process on structural information $\mathbf{c}_{\text{cond}}$ by injecting a residual stream into each block of the denoising model $v_{\theta}$. For a latent representation $\mathbf{z}_{t}$ at a given block, the output $\mathbf{z}^{\prime}_{t}$ is computed as:
@@
$\boldsymbol{F}^{\prime}_{\text{out}}=\text{Attention}({Q}_{\text{tgt}},K^{*},V^{*}).$ (7)
By modulating the injection with our soft mask $M_{S}$, we achieve fine-grained control, preserving non-target regions identified by the TDM while allowing for significant shape modifications in the edit region. This synergy between ControlNet’s geometric enforcement and our TDM-guided semantic preservation enables precise, high-fidelity edits.
Algorithm 1 Region-Controlled Editing
Input: Inference steps $N$, Predicted velocities $\{v_{\text{src}}(\mathbf{x}_{t}^{(i)}),v_{\text{tgt}}(\mathbf{z}_{t}^{(i)})\}_{i=1}^{N}$, Source inversion features $\{K^{\text{inv}}_{t},V^{\text{inv}}_{t}\}_{t=0}^{N}$, target prompt $\mathbf{c}_{\text{tgt}}$, schedule phase durations $\{k_{\text{front}},k_{\text{tail}}\}$
1:for $t=N$ down to 0 do
2:  if $t>k_{\text{front}}$ then
3:   $M_{S}\leftarrow\mathbf{0}$
4:  else if $t>N-k_{\text{front}}-k_{\text{tail}}$ then
5:   $\delta_{t}^{(i)}\leftarrow\left\|v_{\text{tgt}}(\mathbf{z}_{t}^{(i)},t)-v_{\text{src}}(\mathbf{x}_{t}^{(i)},t,\mathbf{c}_{\text{src}})\right\|_{2}$
6:   $\tilde{\delta}_{t}^{(i)}\leftarrow\frac{\delta_{t}^{(i)}-\min_{j}\delta_{t}^{(j)}}{\max_{j}\delta_{t}^{(j)}-\min_{j}\delta_{t}^{(j)}}$
7:   $\hat{\delta}^{(i)}\leftarrow\sum_{t^{\prime}\in T}\frac{\exp(\tilde{\delta}_{t^{\prime}}^{(i)})}{\sum_{t^{\prime\prime}\in T}\exp(\tilde{\delta}_{t^{\prime\prime}}^{(i)})}\cdot\tilde{\delta}_{t^{\prime}}^{(i)}$
8:   $M_{S}\leftarrow\mathcal{G}_{\sigma}*\hat{\delta}$
9:   $M_{S}\leftarrow\mathbf{1}[M_{S}>\tau]$
10:  else
11:   $M_{S}\leftarrow\mathbf{1}$
12:  end if
13:  $K^{*}\leftarrow M_{S}\odot K_{\text{tgt}}+(1-M_{S})\odot K^{\text{inv}}_{t}$
14:  $V^{*}\leftarrow M_{S}\odot V_{\text{tgt}}+(1-M_{S})\odot V^{\text{inv}}_{t}$
15:end for
16:return $K^{*},V^{*}$

### v4

@@ Abstract
footnotetext: † Equal contribution.footnotetext: 🖂 Corresponding author.footnotetext: In Proceedings of the Fourteenth International Conference on Learning Representations (ICLR 2026).
@@ 1 Introduction
Recent advances in generative models have greatly expanded the scope of visual generation, enabling more controllable and realistic modifications across diverse scenarios. Image editing methods based on diffusion [6, 68, 15] and flow models [29, 24, 23, 22] have demonstrated considerable success in general tasks, yet they often fail when faced with complex, large-scale shape transformations. These models can struggle to modify an object’s structure as intended or may inadvertently alter background regions, which degrades the overall image quality. This limitation indicates a critical gap in their ability to perform precise structural edits while maintaining the integrity of unedited content.
@@
To address these challenges, we propose Follow-Your-Shape, a training- and mask-free framework for precise and controllable shape editing. As illustrated in Figure 2, the core innovation of our pipeline is the Trajectory Divergence Map (TDM). The TDM is generated by computing the token-wise difference between the denoising velocity fields of the source and target prompts. This map accurately localizes the regions intended for editing, which in turn guides a selective KV injection mechanism to ensure that modifications are applied precisely where needed while preserving the background.
However, directly applying TDM-guided injection across all denoising timesteps is suboptimal because the TDM can be unstable in the early, high-noise stages of the process. We therefore introduce a Scheduled KV Injection strategy that adapts its guidance throughout the denoising process. As visualized in Figure 2, this staged approach first performs unconditional KV injection to stabilize the initial trajectory, and only then applies TDM-guided editing once a coherent latent structure has formed. This scheduling and staged editing pipeline ensures a more robust and faithful editing outcome compared to a direct application.
To validate our approach, we introduce ReShapeBench, a new benchmark with paired images and refined text prompts specifically designed for evaluating large-scale shape modifications. Beyond this new dataset, we further evaluate Follow-Your-Shape on the public PIE-Bench [20] to ensure generalizability. Follow-Your-Shape achieves state-of-the-art performance on both benchmarks, demonstrating superior background preservation, text–image alignment, and overall visual quality, confirming its effectiveness in both shape-aware and general editing tasks.
@@
  • A novel and training-free editing framework, Follow-Your-Shape, that utilizes a Trajectory Divergence Map (TDM) to achieve precise, large-scale shape transformations while preserving background content.
@@ 2 Related Work
Figure 2: Motivation for Trajectory Divergence Map (TDM) Guided Editing. Left: Vanilla editing methods (red) often produce unstable trajectories compared to the stable reconstruction path (orange). Right: Our staged editing approach better resembles the ideal editing path. The TDM visualizes the dynamically localized editing region across different timesteps, with different border colors corresponding to different stages.
Region-Specific Image Editing.
A central challenge in image editing is localizing modifications to specific regions [4, 90, 30, 18, 39, 41, 33, 62, 12, 15]. Early methods often relied on explicit user-provided masks to delineate editable areas [35, 1, 10, 79, 72]. While effective for certain tasks, this approach requires manual annotation, limiting its applicability. To address this, subsequent work explored techniques to infer editable regions directly from text prompts. Methods such as Prompt-to-Prompt [17] and Plug-and-Play [68] manipulate cross-attention maps to associate textual tokens with spatial areas, enabling localized edits without explicit masks. Other approaches, such as DiffEdit [13], generate a mask by computing differences between diffusion model predictions conditioned on source and target prompts. However, attention-based localization can be imprecise and unstable, especially during large-scale shape transformations where object boundaries change significantly [49, 6]. In contrast, Follow-Your-Shape provides a training-free and mask-free method for identifying editable regions directly from the model’s behavior, avoiding the need for external masks or noisy attention maps.
Structure Preservation via Inversion and Feature Reuse.
Preserving non-target regions is equally critical for high-fidelity editing, and this is closely tied to the quality of the model’s inversion process. For diffusion models, significant research has focused on improving DDIM inversion [63] to better reconstruct a source image from noise. Previous works like null-text inversion [47] and optimization-based methods [70] aim to reduce the discrepancy between the reconstruction and editing trajectories. With the shift toward flow-based models, inversion fidelity has become even more important due to their deterministic nature. RF-Inversion [60] formulates the inversion process as a dynamic optimal control problem, while RF-Solver [73] achieve more accurate reconstructions by incorporating higher-order derivative information. Beyond improving inversion, another line of work focuses on explicitly reusing modules or features from the source image’s generation process [94, 37, 80]. Techniques based on Key-Value (KV) caching [95, 2] or feature injection [73, 15] enforce structural consistency by propagating source-image features into the new generation process. In contrast to prior methods that rely on simple heuristics, Follow-Your-Shape employs a trajectory-guided scheduled injection strategy to achieve more precise, content-aware control.
@@ 3 Methodology
Figure 3: Overview of our proposed pipeline. Given a source image and the corresponding prompt, we first perform inversion to obtain the initial noisy latent code $x_{T}$. The editing process is then divided into three stages. In Stage 1, we stabilize the initial denoising trajectory by reusing key–value (KV) features recorded during inversion and injecting them into the early denoising steps. In Stage 2, we compute a Trajectory Divergence Map (TDM) by comparing denoising trajectories generated under the source and edit prompts, and process this map to precisely localize regions intended for editing. In Stage 3, guided by the TDM, blended KV features are injected into the final attention blocks to introduce new semantics while preserving non-target content. Simultaneously, ControlNet conditions are applied to provide auxiliary structural guidance.
@@ 3.1 Motivation
Effective image editing requires a precise balance between introducing new content and preserving the original structure. As illustrated in Figure 2 (left), traditional structure-preserving editing approaches often produce unstable denoising trajectories that deviate significantly from the stable reconstruction path, leading to severe structural degradation and undesired artifacts. Moreover, prior methods for localizing edits have notable drawbacks:
  • Binary Segmentation Masks: Rely on external tools [21, 59], introducing overhead and a dependency on mask quality. Their rigid boundaries hinder large-scale shape changes and often produce artifacts.
  • Cross-Attention Masks: Inferred from model’s cross attention during the diffusion process, these maps are often noisy and inconsistent, proving unreliable for localizing edits, especially during significant shape transformations.
@@ 3.2 Follow-Your-Shape
We perform shape-aware editing through a staged editing process that combines scheduled Key-Value (KV) injection with structural guidance, where the edit is localized by the Trajectory Divergence Map (TDM).
@@ 3.2.1 Trajectory Divergence Map
Our approach is grounded in the perspective of flow trajectories within the latent space, extending concepts from flow-matching frameworks to the inference setting. As illustrated in Figure 2 (left), a standard reconstruction follows a stable denoising trajectory guided by the source prompt $\mathbf{c}_{\text{src}}$. In an editing task, conditioning on a target prompt $\mathbf{c}_{\text{tgt}}$ alters the velocity field, causing the denoising trajectory to deviate from this initial path. We posit that the magnitude of this deviation spatially localizes the semantic difference between the two prompts. Regions intended for modification will exhibit significant divergence, while background areas will follow nearly identical trajectories. To formalize this, let $\{\mathbf{x}_{t}\}_{t=0}^{T}$ be the latent sequence from the source image inversion, and let $\{\mathbf{z}_{t}\}_{t=0}^{T}$ be the corresponding sequence during the editing (denoising) process. We define the token-wise Trajectory Divergence Map (TDM) $\delta_{t}$ at timestep $t$ as the $L_{2}$ norm of the difference between the velocity vectors predicted under the two prompts:
@@
As shown in Figure 2 (right), this produces a normalized TDM, $\{\tilde{\delta}_{t}^{(i)}\}$, which quantifies the localized editing strength on a scale of $[0,1]$.
3.2.2 Staged Editing and Structural Guidance
Directly applying TDM-guided injection across all timesteps is suboptimal due to the instability of the TDM in early, high-noise regimes (Figure 2 right). Early latents provide weak and noisy spatial signals, which can mislocalize edits if aggressive guidance is applied too soon. To address this, we introduce a scheduled injection strategy that partitions the $N$ denoising steps into three distinct phases and adapts the guidance mechanism to the latent state: the first phase emphasizes stabilization, the second collects and aggregates TDM evidence while allowing exploration, and the third enforces structural and semantic conformance.
@@ 3.2.2 Staged Editing and Structural Guidance
For an initial set of $k_{\text{front}}$ timesteps, we perform unconditional KV injection from the source inversion path across all spatial tokens. This operation enforces a global reconstruction objective, equivalent to setting the edit mask $M_{S}=\mathbf{0}$, which stabilizes the trajectory and prevents semantic drift while the latent representation $\mathbf{z}_{t}$ is still dominated by noise. Intuitively, the model first anchors to a faithful reconstruction manifold before any region-specific modification is attempted, reducing the risk of spurious changes to background layout or texture.
Stage 2: Editing and TDM Aggregation.
Once a stable latent structure has emerged, we begin the editing phase over a predefined window of timesteps $N$. During this window, we perform editing by setting the edit mask $M_{S}=1$ at every step, allowing the model to explore target-guided generation path. Simultaneously, we compute and store the normalized TDMs $\tilde{\delta}_{t}$ at each timestep within $N$, capturing the trajectory divergence guided by the source and target prompts. After this editing window concludes, we aggregate the stored TDMs $\{\tilde{\delta}_{t}\}$ across time to construct a temporally consistent and spatially coherent edit mask. Specifically, throughout the denoising process, a token that appears unchanged at an individual timestep may still experience evolution at subsequent steps. Therefore, to ensure that the aggregation faithfully captures such temporal dynamics, we employ a softmax-weighted temporal fusion for each token $i$:
$\hat{\delta}^{(i)}=\sum_{t\in N}\alpha_{t}^{(i)}\cdot\tilde{\delta}_{t}^{(i)},\quad\text{where}\quad\alpha_{t}^{(i)}=\frac{\exp(\tilde{\delta}_{t}^{(i)})}{\sum_{t^{\prime}\in N}\exp(\tilde{\delta}_{t^{\prime}}^{(i)})}.$ (3)
To ensure spatial coherence and suppress noisy edges, the resulting map $\hat{\delta}$ is further refined via convolution with a Gaussian kernel $\mathcal{G}_{\sigma}$ to obtain $\tilde{M}_{S}\in[0,1]^{H\times W}$:
$\tilde{M}_{S}=\mathcal{G}_{\sigma}*\hat{\delta}$. (4)
We observe that the distribution of values in $\tilde{M}_{S}$ typically exhibits a skewed unimodal shape (as shown in Figure 4), characterized by a dominant background mode and a long-tailed foreground response. Such a distribution is well suited for Otsu’s method [48], which selects the threshold $\tau$ that that maximizes the inter-class variance of values. Formally, for a candidate threshold $\tau$, let $\omega_{0},\mu_{0}$ and $\omega_{1},\mu_{1}$ denote the class probabilities and means of the background ($\tilde{M}_{S}\leq\tau$) and foreground ($\tilde{M}_{S}>\tau$), respectively. The between-class variance is defined as:
$\sigma_{b}^{2}(\tau)=\omega_{0}(\tau)\,\omega_{1}(\tau)\,\big(\mu_{0}(\tau)-\mu_{1}(\tau)\big)^{2}.$ (5)
Figure 4: Distribution of values of $\tilde{M}_{S}$. The red dashed line indicates the Otsu threshold $\tau$.
The optimal threshold is then given by:
$\tau^{*}=\arg\max_{\tau}\,\sigma_{b}^{2}(\tau).$ (6)
The final binary mask $M_{S}$ is thus obtained by applying this threshold:
$M_{S}=\mathbf{1}\!\left[\tilde{M}_{S}>\tau^{*}\right],\quad M_{S}\in\{0,1\}^{H\times W},$ (7)
where $\mathbf{1}[\cdot]$ denotes the indicator function.
Figure 5: Construction Process of ReShapeBench. Note that images in Step 3 are generated after benchmark construction to serve as visual references. Checklist validation is performed on prompt.
@@
Our framework enforces structural conformance by jointly leveraging TDM-guided feature injection for background preservation and ControlNet residual conditioning for stabilizing structural patterns. The mask $M_{S}$ obtained in Stage 2 modulates the fusion of Key-Value features, activating the target features ($K_{\text{tgt}},V_{\text{tgt}}$) in edited regions and reverting to the source features ($K^{\text{inv}},V^{\text{inv}}$) elsewhere. This feature-blending operation is formulated as:
$\{K^{*},V^{*}\}\leftarrow M_{S}\odot\{K^{\text{tgt}},V^{\text{tgt}}\}+(1-M_{S})\odot\{K^{\text{inv}},V^{\text{inv}}\}.$ (8)
For structural guidance, ControlNet conditions the process on structural information $\mathbf{c}_{\text{cond}}$ by injecting a residual stream into each block of the denoising model $v_{\theta}$. For a latent representation $\mathbf{z}_{t}$ at a given block, the output $\mathbf{z}^{\prime}_{t}$ is computed as:
@@
$\boldsymbol{F}^{\prime}_{\text{out}}=\text{Attention}({Q}^{\text{tgt}},K^{*},V^{*}).$ (10)
This synergy between ControlNet’s geometric enforcement and our TDM-guided semantic preservation enables precise, high-fidelity edits. The algorithmic implementation can be found in Algorithm 1 in Supplementary Material.
4 ReShapeBench: A Benchmark for Large-Scale Shape Transformations
Overview.
Existing benchmarks for image editing [77, 20, 86] are not tailored to the demands of shape-aware editing, where the goal is to change object geometry while preserving the surrounding background. In particular, PIE-Bench [20] (700 images) uses concise prompts that often lack spatial or structural detail, and it aggregates heterogeneous tasks (object replacement, stylization, background modification) rather than isolating shape transformation as a first-class target. These properties make it difficult to diagnose whether a method truly performs structural change or relies on side effects such as texture shifts or background re-synthesis. We therefore introduce ReShapeBench, a benchmark that centers on mask-free, prompt-driven shape transformation with paired prompts and controlled background settings. This design isolates the factors relevant to structural change and reduces confounds from style or background alterations, enabling a targeted and reproducible evaluation protocol. It also serves as a targeted complement to existing evaluation suites, providing shape-focused test cases that fill the gap left by current general-purpose editing benchmarks.
Benchmark Construction.
ReShapeBench contains 120 newly collected images split into three subsets: 70 single-object scenes for precise shape editing, 50 multi-object scenes for targeted mask-free edits, and a general evaluation set of 50 images that combines samples from both subsets with curated PIE-Bench cases to assess generalization. All images are standardized to $512\times 512$ to normalize spatial scale and reduce variability across methods and backbones. Each new image is paired with two distinct shape transformations, yielding 240 editing cases across the single- and multi-object subsets, plus 50 cases in the general set, for a total of 290 shape-aware editing cases; this pairing increases task coverage and controls difficulty by varying the magnitude of structural change. Source–target prompts follow a structured template and differ only in the foreground object description, which stabilizes text-to-image alignment while holding background fixed. All prompt pairs are generated by Qwen-2.5-VL [3] and validated by human raters to ensure alignment and that the transformation satisfies the predefined shape criteria; ambiguous cases are double-checked to maintain consistency. Figure 5 and the Supplementary Material 8 present the construction procedure and sample cases, including the selection checklist and prompt schema used during curation.